OpenStudy (anonymous):

Find two positive numbers A and B (with A≤B) whose sum is 44 and whose product is maximized

OpenStudy (anonymous):

so A+B=44 now let P be the product of the two so P=AB notice that using the first equation, B=44-A, so the product becomes P=A(44-A) P=44A-A^2 now this is just a single variable calc. problem about optimization. So we find the critical points(set P'=0 and find the A there). P'=44-2A P'=0 44-2A=0 -2A=-44 A=22 since P''(22)<0 then we know A=22 is a maximum.. so the first number is x=22 and the second number will be B=44-A=44-22=22 So A=22 and B=22 will yield the maximum product

OpenStudy (anonymous):

sorry typographical error.. change the x at the line " so the first number is x=22 and the second number will be B=44-A=44-22=22" to A

OpenStudy (anonymous):

got used to x :))

OpenStudy (anonymous):

Find two numbers A and B (with A≤B) whose difference is 38 and whose product is minimized if it had asked me to do minimized

OpenStudy (anonymous):

ok notice that this will be a single var. calc optimization prob again. Here we have B-A=38 let P be the product of the two P=AB we know that P=A(A+38) P=A^2+38A find the critical points again: P'=2A+38 P'=0 2A+38=0 2A=-38 A=-19 this will be the minimum since P''(-19)>0... so B will be: B=38+A=38-19=19 the answer will be B=19 and A=-19.